Violet knows that her subscription rate for cable increased from $18.20 to $19.50 at some point during the calendar year. She paid a total of $230.10 for the year to the cable company. Determine the month of the rate increase.
Answer by ankor@dixie-net.com(22740)

Violet knows that her subscription rate for cable increased from $18.20 to $19.50
at some point during the calendar year.
She paid a total of $230.10 for the year to the cable company.
Determine the month of the rate increase.
:
let x = the no. months at 18.20
then
(12-x) = no. of months at 19.50
:
18.20x + 19.50(12-x) = 230.10
18.20x + 234 - 19.50x = 230.10
18.20x - 19.50x = 230.10 - 234
-1.3x = -3.9
x = -3.9/-1.3
x = +3 months at the original rate
The new rate came in the 4th month, April
:
:
See if that checks out
3(18.20) = 54.60
9(19.50) = 175.50
-------------------
total for yr: 230.10